OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 2) = 12

OpenStudy (shane_b):

Do you know where to start?

OpenStudy (anonymous):

I believe I would do the Distributive Property first for 4(x-1) Is this correct?

OpenStudy (shane_b):

Yes...so you get?

OpenStudy (anonymous):

\[4(x)-4(1) \] = \[4x-4-(x+2)=12\] and then ?

OpenStudy (shane_b):

You're doing well so far...what do you think happens next?

OpenStudy (shane_b):

(hint: the - sign on the far right is essentially a -1)

OpenStudy (anonymous):

I believe then I would do Distributive Property for 4(x+2) ?

OpenStudy (shane_b):

No. You've already removed the parentheses on the left through distribution...you need to do the same thing with -(x+2)....get rid of those parentheses!

OpenStudy (anonymous):

so 4(-x+2)?

OpenStudy (shane_b):

Wow....we were getting there @mathslover

mathslover (mathslover):

OOps sorry !! for disturbing u

OpenStudy (anonymous):

so would it be 4(-x+2) for Distribution?

OpenStudy (shane_b):

@mathslover : thx :)

OpenStudy (shane_b):

@verdoc: No....we're only working with -(x+2) at the moment. Can you distribute the negative sign there and remove the parentheses?

OpenStudy (anonymous):

so then it would be -x+-2?

OpenStudy (shane_b):

Close enough...properly written it will be -x-2. Ok, we're almost done. If you put everything back together you get 4x-4-x-2=12 right?

OpenStudy (anonymous):

Yes. So then combine like terms and get 5x -4-2=12?

OpenStudy (shane_b):

Yes combine like terms...but correctly :) Can you see any mistakes there?

OpenStudy (anonymous):

Subtract!! 3x-4-2=12!

OpenStudy (shane_b):

Now what's -4-2? Need to simplify that too

OpenStudy (anonymous):

That equals -6.

OpenStudy (shane_b):

Correct. So you have 3x-6=12. What's next?

OpenStudy (anonymous):

Add -6 to 12.?

OpenStudy (shane_b):

Right...add +6 to both sides. That leaves you with what?

OpenStudy (anonymous):

3x=18 then divide both sides by 3 and get x= 6!

OpenStudy (shane_b):

\[\large BINGO\]

OpenStudy (anonymous):

YESSSSS!!!!!!! thank you!!

OpenStudy (shane_b):

you're welcome. With a little practice you can do this all in your head in just a few seconds :)